3x-y+2z=17
4x+3y+5z=12
6x+8y+4z=4

Answer by richwmiller(17219)

3,-1,2,17
4,3,5,12
6,8,4,4

determinant = -70
x = -280/-70 = 4/1 = 4
y = 210/-70 = -3/1 = -3
z = -70/-70 = 1/1 = 1

1,0,0,4
0,1,0,-3
0,0,1,1